OpenStudy (anonymous):

Evaluate the line integral zdx+xdy+xydz where is the path of the helix r(t)=(2cost)i+(2sint)j+(1t)k on 0 less than t less than 2

OpenStudy (anonymous):

make variable change to cyulindrical coordinates first with r=2

OpenStudy (anonymous):

x=2cost y=2sint z=z so: dx=-2sint*dt dy=2cost*dt dz=dz

OpenStudy (anonymous):

i got to \[\int\limits_{0}^{2\pi} -2 tsintdt +4\cos^2tdt + 4costsintdt\] but i cant figure put how to integrate

OpenStudy (turingtest):

t is bound above by 2 or 2pi?

OpenStudy (turingtest):

first term: integrate by parts second term: double angle formula \(\cos^2t=\frac12(1+\cos(2t))\) third term: u=sin(t)

OpenStudy (anonymous):

\[0\le t \le 2\pi\]

OpenStudy (turingtest):

that makes more sense obviously so did you try it yet?

OpenStudy (anonymous):

no the homework is not due till tonight so im studying for my test and will try it after!

OpenStudy (turingtest):

good luck!

OpenStudy (anonymous):

i got 3 pi but it was incorrect.

OpenStudy (turingtest):

\[\int_0^{2\pi}-2t\sin t+4\cos^2t+4\sin t\cos tdt\]\[2t\cos t-2\sin t+2t+\sin(2t)+2\sin t|_0^{2\pi}\]is that what you got?

OpenStudy (anonymous):

no my signs were messed up.

OpenStudy (turingtest):

let me check with wolfram...

OpenStudy (anonymous):

thats correct! i finished it!

OpenStudy (anonymous):

i had too many subtraction signs.

OpenStudy (turingtest):

cool :) yeah, the algebra can be the worst in calc 3

OpenStudy (anonymous):

hahhaa. yeah. thanks so much. :)